OpenStudy (anonymous):

find the equastion in slope intercept form for the line contaning the two points show work. (: ( 3,12 ) and (-4,-16 )

OpenStudy (anonymous):

equation*

OpenStudy (anonymous):

that is the equastion/

OpenStudy (anonymous):

find slope and then solve for b

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

follow along :P

OpenStudy (anonymous):

kayy gooo

OpenStudy (anonymous):

y2 - y1 m = ------ x2 - x1 -16 - 12 m = ------- -4 - 3 -28 m = -- -7 m = 4

OpenStudy (anonymous):

thats slope

OpenStudy (anonymous):

y = mx+b use any coordinate 12 = 4(3) +b 12 = 12 + b 12 - 12 = b 0 = b y = 4x

OpenStudy (anonymous):

So that would be the equastion ^ ^

OpenStudy (anonymous):

yeah :)

OpenStudy (anonymous):

thanks okay can you show me how to do this one. but step by step (-3,7 ) and ( 6,-8 )

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

wait here do this one (-3,7) and (6,-8)

OpenStudy (anonymous):

sorry my computer crashed a while ago and sure ill do it

OpenStudy (anonymous):

ok look i want you to learn alrii soo imma tell you what to do and you do it step by step :P

OpenStudy (anonymous):

call me your tutor ;)

OpenStudy (anonymous):

ok so first off you have your two points substitute the values into the equation for slope

OpenStudy (anonymous):

okay tutor thanks your the best, (: okay what does subsitute mean .?

OpenStudy (anonymous):